# Beyond the Linear Separability Ceiling

### v1

@@ Appendix H PEFT comparison
LoRA ($\mathcal{L}_{\text{NT}}$) $\underline{\mathbf{77.4\pm 2.9}}$ $\mathbf{71.9\pm 3.1}$ $71.1\pm 3.1^{I,O}$ $61.1\pm 3.4^{I}$ $63.1\pm 3.3$
LoRA ($\mathcal{L}_{\text{combined}}$) $\underline{\mathbf{78.9\pm 2.8}}$ $\mathbf{71.2\pm 3.1}$ $71.8\pm 3.1^{I,O}$ $79.4\pm 2.8^{I}$ $79.4\pm 2.8^{O}$
Pixtral OpenWorld Direct baseline $\mathbf{72.4\pm 3.9}$ $\mathbf{79.4\pm 3.5}$ $76.0\pm 3.7^{I,O}$ $87.2\pm 2.9^{I}$ $88.0\pm 2.8^{O}$
Postfix tuning ($\mathcal{L}_{\text{NT}}$) $\underline{\mathbf{93.6\pm 2.1}}$ $\underline{\mathbf{93.6\pm 2.1}}$ $76.6\pm 3.7$ $87.6\pm 2.9$ $87.6\pm 2.9$
Prompt tuning ($\mathcal{L}_{\text{NT}}$) $\underline{\mathbf{93.6\pm 2.1}}$ $\mathbf{79.2\pm 3.8}$ $76.6\pm 3.7^{I}$ $78.4\pm 3.6$ $83.0\pm 3.3^{O}$
Prompt tuning ($\mathcal{L}_{\text{combined}}$) $\underline{\mathbf{94.4\pm 2.0}}$ $\underline{\mathbf{88.2\pm 2.8}}$ $76.6\pm 3.7$ $95.0\pm 1.9^{I}$ $93.4\pm 2.2^{O}$
LoRA ($\mathcal{L}_{\text{NT}}$) $\underline{\mathbf{93.4\pm 2.2}}$ $\underline{\mathbf{90.0\pm 2.6}}$ $76.6\pm 3.7$ $87.2\pm 2.9^{I}$ $89.6\pm 2.7$
LoRA ($\mathcal{L}_{\text{combined}}$) $\underline{\mathbf{95.0\pm 1.9}}$ $\underline{\mathbf{94.6\pm 2.0}}$ $74.4\pm 3.8$ $96.2\pm 1.7^{I}$ $95.4\pm 1.8^{O}$
HOI Direct baseline $\mathbf{57.8\pm 3.4}$ $\mathbf{60.6\pm 3.4}$ $62.7\pm 3.4^{I,O}$ $70.2\pm 3.2^{I}$ $70.5\pm 3.2$
Postfix tuning ($\mathcal{L}_{\text{NT}}$) $\mathbf{66.4\pm 3.3}$ $\mathbf{66.4\pm 3.3}$ $62.7\pm 3.4^{I,O}$ $71.0\pm 3.1^{I}$ $71.0\pm 3.1^{O}$
Prompt tuning ($\mathcal{L}_{\text{NT}}$) $\mathbf{58.6\pm 3.4}$ $54.0\pm 3.5$ $62.7\pm 3.4^{I}$ $62.6\pm 3.4^{I}$ $63.1\pm 3.3^{-O}$
Prompt tuning ($\mathcal{L}_{\text{combined}}$) $\underline{\mathbf{70.8\pm 3.2}}$ $\mathbf{58.9\pm 3.4}$ $62.7\pm 3.4^{I}$ $73.5\pm 3.1^{I}$ $72.2\pm 3.1^{O}$
LoRA ($\mathcal{L}_{\text{NT}}$) $\underline{\mathbf{78.0\pm 2.9}}$ $\underline{\mathbf{74.4\pm 3.0}}$ $61.6\pm 3.4^{I,O}$ $74.9\pm 3.0^{I}$ $73.2\pm 3.1^{O}$
LoRA ($\mathcal{L}_{\text{combined}}$) $\underline{\mathbf{79.6\pm 2.8}}$ $\mathbf{62.0\pm 3.4}$ $63.1\pm 3.3^{I}$ $77.8\pm 2.9^{I}$ $77.6\pm 2.9^{O}$
Gemma3 4B OpenWorld Direct baseline $76.0\pm 3.7$ $68.6\pm 4.1$ $89.8\pm 2.7$ $50.0\pm 4.4^{I}$ $50.0\pm 4.4^{O}$
Postfix tuning ($\mathcal{L}_{\text{NT}}$) $\mathbf{86.0\pm 3.0}$ $\mathbf{86.0\pm 3.0}$ $89.8\pm 2.7$ $50.0\pm 4.4$ $50.0\pm 4.4$
Prompt tuning ($\mathcal{L}_{\text{NT}}$) $\mathbf{86.0\pm 3.0}$ $71.2\pm 4.0$ $89.8\pm 2.7^{I}$ $50.0\pm 4.4$ $50.2\pm 4.4^{O}$
@@
LoRA ($\mathcal{L}_{\text{NT}}$) $\mathbf{92.4\pm 2.3}$ $\mathbf{94.0\pm 2.1}$ $89.8\pm 2.7^{I}$ $50.0\pm 4.4$ $50.0\pm 4.4$
LoRA ($\mathcal{L}_{\text{combined}}$) $\underline{\mathbf{95.6\pm 1.8}}$ $\underline{\mathbf{96.2\pm 1.7}}$ $89.8\pm 2.7^{O}$ $96.6\pm 1.6^{I}$ $96.8\pm 1.5^{O}$
HOI Direct baseline $56.5\pm 3.4$ $54.2\pm 3.5$ $74.1\pm 3.0$ $50.0\pm 3.5^{I}$ $50.0\pm 3.5^{O}$
Postfix tuning ($\mathcal{L}_{\text{NT}}$) $56.1\pm 3.4$ $56.1\pm 3.4$ $74.1\pm 3.0$ $50.0\pm 3.5^{-I}$ $50.0\pm 3.5^{-O}$
Prompt tuning ($\mathcal{L}_{\text{NT}}$) $54.1\pm 3.5$ $51.6\pm 3.5$ $74.1\pm 3.0^{O}$ $50.0\pm 3.5^{I}$ $50.0\pm 3.5^{-O}$
Prompt tuning ($\mathcal{L}_{\text{combined}}$) $58.9\pm 3.4$ $54.0\pm 3.5$ $74.1\pm 3.0^{I}$ $50.0\pm 3.5^{-I}$ $50.0\pm 3.5^{O}$
LoRA ($\mathcal{L}_{\text{NT}}$) $\underline{\mathbf{84.2\pm 2.5}}$ $\underline{\mathbf{81.8\pm 2.7}}$ $74.1\pm 3.0^{I,O}$ $50.0\pm 3.5$ $50.0\pm 3.5$
LoRA ($\mathcal{L}_{\text{combined}}$) $\underline{\mathbf{84.2\pm 2.5}}$ $67.4\pm 3.2$ $74.1\pm 3.0^{I}$ $83.2\pm 2.6^{I}$ $82.1\pm 2.7^{O}$
Table 7 presents results from these explorations. Key finding here is that the performance gains achieved by the $\mathcal{L}_{\text{combined}}$ objective for prompt tuning are, for the most part, not unique to that loss function. As shown, the “fast” configuration achieves generative accuracy that is highly comparable to the main paper’s prompt tuning ($\mathcal{L}_{\text{combined}}$) baseline. This suggests that multiple pathways exist to surpass the LSC. While the $\mathcal{L}_{\text{combined}}$ objective provides a more robust method for aligning the embedding space, aggressive hyperparameter tuning with a standard language modeling objective can be similarly effective at unlocking a model’s latent reasoning capabilities.
Also, LoRAs trained with the $\mathcal{L}_{\text{NT}}$ objective demonstrate stronger structural generalization. Across all models and datasets, these versions largely maintain their performance when evaluated on the OOD prompt format, showing no statistically significant degradation. In contrast, LoRAs trained with the $\mathcal{L}_{\text{combined}}$ objective sometimes exhibit structural brittleness. In half of the tested scenarios (specifically, all models on the HOI dataset), their generative performance drops significantly on the OOD prompt structure, even while the corresponding final-layer similarity evaluation remains high. A similar, and even clearer, pattern emerges when comparing prompt-based tuning methods. Postfix tuning, which only appends learnable tokens, proves to be highly robust, maintaining its performance on OOD formats. In contrast, prompt tuning, which includes a learnable prefix, is more susceptible to syntactic overfitting and shows a more significant performance drop on OOD prompts. This reinforces that interventions at the start of the sequence are more likely to create a dependency on prompt syntax than those that only modify the end. This suggests that while explicitly optimizing for representation quality is effective, it can inadvertently encourage the generative model to rely on structural shortcuts, harming its ability to generalize to new prompt formats.
Appendix I Results across HOI splits
Table 8 provides a detailed breakdown of model performance across the four distinct test splits of the Bongard-HOI dataset. These results are benchmarked against two state-of-the-art methods: Test-Time Prompt Tuning (TPT) (Shu et al., 2022), which adapts prompts for vision-language models at test time, and SVM-Mimic (Raghuraman et al., 2024), a Transformer-based approach that learns from SVM-derived support sets.
Table 8: Performance on HOI - detailed breakdown
@@ Appendix I Results across HOI splits
Avg acc. (%)
Phi Baseline gen. 51.5 55.0 51.0 51.0 52.1
Baseline sim. 61.5 60.0 57.5 63.0 60.5
Prompt tuning gen. 60.5 59.0 52.5 62.0 58.5
Prompt tuning sim. 61.5 59.0 57.0 62.0 59.9
Prompt tuning gen. $\mathcal{L}_{\text{sim}}$ 66.5 69.0 62.5 63.5 65.4
Prompt tuning sim. $\mathcal{L}_{\text{sim}}$ 63.5 73.5 70.5 67.5 68.8
LoRA gen. $\mathcal{L}_{\text{NT}}$ 78.0 77.5 75.0 79.0 77.4
LoRA sim. $\mathcal{L}_{\text{NT}}$ 63.5 59.0 60.0 62.0 61.1
LoRA gen. $\mathcal{L}_{\text{combined}}$ 78.0 79.5 75.5 82.5 78.9
LoRA sim. $\mathcal{L}_{\text{combined}}$ 80.5 77.0 76.5 83.5 79.4
Pixtral Baseline gen. 60.0 61.5 51.5 58.0 57.8
Baseline sim. 66.5 73.5 69.5 71.5 70.2
Prompt tuning gen. 60.5 58.5 56.5 59.0 58.6
Prompt tuning sim. 63.0 62.0 62.5 63.0 62.6
Prompt tuning gen. $\mathcal{L}_{\text{sim}}$ 73.0 74.0 64.0 72.0 70.8
Prompt tuning sim. $\mathcal{L}_{\text{sim}}$ 76.0 73.5 70.0 74.5 73.5
LoRA gen. $\mathcal{L}_{\text{NT}}$ 77.0 79.5 75.5 80.0 78.0
LoRA sim. $\mathcal{L}_{\text{NT}}$ 73.0 77.5 71.0 78.0 74.9
LoRA gen. $\mathcal{L}_{\text{combined}}$ 78.5 81.5 74.5 84.0 79.6
LoRA sim. $\mathcal{L}_{\text{combined}}$ 79.5 78.5 74.5 78.5 77.8
Gemma3 4B Baseline gen. 58.5 56.0 53.0 58.5 56.5
Baseline sim. 50.0 50.0 50.0 50.0 50.0
Prompt tuning gen. 53.0 56.5 53.0 54.0 54.1
Prompt tuning sim. 50.0 50.0 50.0 50.0 50.0
Prompt tuning gen. $\mathcal{L}_{\text{sim}}$ 56.5 61.5 59.5 58.0 58.9
Prompt tuning sim. $\mathcal{L}_{\text{sim}}$ 50.0 50.0 50.0 50.0 50.0
LoRA gen. $\mathcal{L}_{\text{NT}}$ 83.5 84.5 85.0 84.0 84.2
LoRA sim. $\mathcal{L}_{\text{NT}}$ 50.0 50.0 50.0 50.0 50.0
LoRA gen. $\mathcal{L}_{\text{combined}}$ 83.0 84.0 85.0 85.0 84.2
LoRA sim. $\mathcal{L}_{\text{combined}}$ 84.0 86.5 82.5 80.0 83.2
CLIP-RN50 TPT 66.4 68.5 66.0 65.5 66.6
SVM-Mimic 69.6 70.8 78.1 71.2 72.5
@@
All three models fine-tuned with LoRA significantly outperform not only their own baselines but also the previous state-of-the-art methods, TPT and SVM-Mimic. The most striking finding is the consistent high performance across all splits. There is no significant accuracy degradation when models are tested on unseen objects, unseen actions, or both simultaneously. In fact, for both Phi and Pixtral, the highest generative accuracy with the combined loss is achieved on the most difficult split (unseen obj. / unseen act.), providing strong evidence that the models are learning the abstract relational concept rather than memorizing training examples.
Appendix J Domain generalization
@@
In this evaluation, models were fine-tuned on one dataset (e.g., OpenWorld) and evaluated on the other (e.g., HOI) to assess their ability to generalize beyond the training domain’s specific concepts and structures. The dataset column in the table indicates the evaluation dataset, where the evaluation is done on the other dataset.
@@
Table 9: Cross-domain generalization performance. Reported accuracies presented as $accuracy\pm ME$. bold indicates surpassing of LSC. Superscripts on similarity scores denote a significant dependence (p < 0.05) with the predictions of the corresponding generative method (G).
@@
LoRA ($\mathcal{L}_{\text{NT}}$) $\mathbf{92.6\pm 2.3}$ $84.2\pm 3.2^{G}$ $76.0\pm 3.7^{G}$
LoRA ($\mathcal{L}_{\text{combined}}$) $\mathbf{93.8\pm 2.1}$ $84.2\pm 3.2^{G}$ $87.8\pm 2.9^{G}$
@@
LoRA ($\mathcal{L}_{\text{NT}}$) $62.1\pm 3.4$ $71.1\pm 3.1^{G}$ $59.5\pm 3.4^{-G}$
LoRA ($\mathcal{L}_{\text{combined}}$) $67.6\pm 3.2$ $71.8\pm 3.1^{G}$ $66.9\pm 3.3^{G}$
@@
LoRA ($\mathcal{L}_{\text{NT}}$) $\mathbf{88.8\pm 2.8}$ $76.6\pm 3.7^{G}$ $89.0\pm 2.7^{G}$
LoRA ($\mathcal{L}_{\text{combined}}$) $\mathbf{83.4\pm 3.3}$ $74.4\pm 3.8^{G}$ $87.4\pm 2.9^{G}$
@@
LoRA ($\mathcal{L}_{\text{NT}}$) $64.9\pm 3.3$ $61.6\pm 3.4^{G}$ $72.5\pm 3.1^{G}$
LoRA ($\mathcal{L}_{\text{combined}}$) $\mathbf{71.0\pm 3.1}$ $63.1\pm 3.3^{G}$ $68.4\pm 3.2^{G}$
@@
LoRA ($\mathcal{L}_{\text{NT}}$) $87.2\pm 2.9$ $89.8\pm 2.7^{G}$ $50.0$
LoRA ($\mathcal{L}_{\text{combined}}$) $83.4\pm 3.3$ $89.8\pm 2.7$ $83.2\pm 3.3^{G}$

### v3

@@ Appendix I Combined loss hyperparameter search
LoRA ($\mathcal{L}_{\text{NT}}$) $\mathbf{78.6\pm 2.8}$ $72.8\pm 3.1$ $71.9\pm 3.1^{I,O}$ $63.6\pm 3.3^{I}$ $64.4\pm 3.3$
LoRA ($\mathcal{L}_{\text{combined}}$) $\mathbf{79.2\pm 2.8}$ $73.5\pm 3.1$ $71.9\pm 3.1^{I,O}$ $82.0\pm 2.7^{I}$ $82.9\pm 2.6^{O}$
Pixtral OpenWorld Direct baseline $72.4\pm 3.9$ $79.4\pm 3.5$ $76.0\pm 3.7^{I,O}$ $7.2\pm 2.9^{I} 88$$8.0\pm 2.8^{O}$
Postfix tuning ($\mathcal{L}_{\text{NT}}$) $\mathbf{93.6\pm 2.1}$ $\mathbf{93.6\pm 2.1}$ $76.6\pm 3.7$ $87.6\pm 2.9$ $87.6\pm 2.9$
Prompt tuning ($\mathcal{L}_{\text{NT}}$) $\mathbf{93.6\pm 2.1}$ $79.2\pm 3.8$ $76.6\pm 3.7^{I}$ $78.4\pm 3.6$ $83.0\pm 3.3^{O}$
Prompt tuning ($\mathcal{L}_{\text{combined}}$) $\mathbf{94.4\pm 2.0}$ $\mathbf{88.2\pm 2.8}$ $76.6\pm 3.7$ $95.0\pm 1.9^{I}$ $93.4\pm 2.2^{O}$
LoRA ($\mathcal{L}_{\text{NT}}$) $\mathbf{93.4\pm 2.2}$ $\mathbf{90.0\pm 2.6}$ $76.6\pm 3.7$ $87.2\pm 2.9^{I}$ $89.6\pm 2.7$
LoRA ($\mathcal{L}_{\text{combined}}$) $\mathbf{95.0\pm 1.9}$ $\mathbf{94.6\pm 2.0}$ $74.4\pm 3.8$ $96.2\pm 1.7^{I}$ $95.4\pm 1.8^{O}$
HOI Direct baseline $57.8\pm 3.4$ $60.6\pm 3.4$ $62.7\pm 3.4^{I,O}$ $70.2\pm 3.2^{I}$ $70.5\pm 3.2$
Postfix tuning ($\mathcal{L}_{\text{NT}}$) $6.4\pm 3.3 66$$6.4\pm 3.3$ $62.7\pm 3.4^{I,O}$ $71.0\pm 3.1^{I}$ $71.0\pm 3.1^{O}$
Prompt tuning ($\mathcal{L}_{\text{NT}}$) $58.6\pm 3.4$ $54.0\pm 3.5$ $62.7\pm 3.4^{I}$ $62.6\pm 3.4^{I}$ $63.1\pm 3.3^{-O}$
Prompt tuning ($\mathcal{L}_{\text{combined}}$) $\mathbf{70.8\pm 3.2}$ $58.9\pm 3.4$ $62.7\pm 3.4^{I}$ $73.5\pm 3.1^{I}$ $72.2\pm 3.1^{O}$
LoRA ($\mathcal{L}_{\text{NT}}$) $\mathbf{78.0\pm 2.9}$ $\mathbf{74.4\pm 3.0}$ $61.6\pm 3.4^{I,O}$ $74.9\pm 3.0^{I}$ $73.2\pm 3.1^{O}$
LoRA ($\mathcal{L}_{\text{combined}}$) $\mathbf{79.6\pm 2.8}$ $62.0\pm 3.4$ $63.1\pm 3.3^{I}$ $7.8\pm 2.9^{I} 77$$7.6\pm 2.9^{O}$
Gemma3 4B OpenWorld Direct baseline $76.0\pm 3.7$ $68.6\pm 4.1$ $89.8\pm 2.7$ $50.0^{I}$ $50.0^{O}$
Postfix tuning ($\mathcal{L}_{\text{NT}}$) $86.0\pm 3.0$ $86.0\pm 3.0$ $89.8\pm 2.7$ $50.0$ $50.0$
Prompt tuning ($\mathcal{L}_{\text{NT}}$) $86.0\pm 3.0$ $71.2\pm 4.0$ $89.8\pm 2.7^{I}$ $50.0$ $50.2\pm 4.4^{O}$
@@
LoRA ($\mathcal{L}_{\text{NT}}$) $92.4\pm 2.3$ $\mathbf{94.0\pm 2.1}$ $89.8\pm 2.7^{I}$ $50.0$ $50.0$
LoRA ($\mathcal{L}_{\text{combined}}$) $\mathbf{95.6\pm 1.8}$ $\mathbf{96.2\pm 1.7}$ $89.8\pm 2.7^{O}$ $96.6\pm 1.6^{I}$ $96.8\pm 1.5^{O}$
HOI Direct baseline $56.5\pm 3.4$ $54.2\pm 3.5$ $74.1\pm 3.0$ $50.0^{I}$ $50.0^{O}$
Postfix tuning ($\mathcal{L}_{\text{NT}}$) $56.1\pm 3.4$ $56.1\pm 3.4$ $74.1\pm 3.0$ $50.0^{-I}$ $50.0^{-O}$
Prompt tuning ($\mathcal{L}_{\text{NT}}$) $54.1\pm 3.5$ $51.6\pm 3.5$ $74.1\pm 3.0^{O}$ $50.0^{I}$ $50.0^{-O}$
Prompt tuning ($\mathcal{L}_{\text{combined}}$) $58.9\pm 3.4$ $54.0\pm 3.5$ $74.1\pm 3.0^{I}$ $50.0^{-I}$ $50.0^{O}$
LoRA ($\mathcal{L}_{\text{NT}}$) $\mathbf{84.2\pm 2.5}$ $\mathbf{81.8\pm 2.7}$ $74.1\pm 3.0^{I,O}$ $50.0$ $50.0$
LoRA ($\mathcal{L}_{\text{combined}}$) $\mathbf{84.2\pm 2.5}$ $67.4\pm 3.2$ $74.1\pm 3.0^{I}$ $83.2\pm 2.6^{I}$ $82.1\pm 2.7^{O}$
The results in Table A.5 reveal two key insights regarding the methods for resolving the alignment gap for this task.
Mechanistic divergence towards a common goal. A primary finding is that multiple strategies exist to achieve high generative accuracy. For instance, postfix tuning with aggressive hyperparameters can yield generative performance comparable to methods using the $\mathcal{L}_{\text{combined}}$ objective. However, this equivalence in outcome should not be mistaken for an equivalence in mechanism. The aggressive $\mathcal{L}_{\text{NT}}$ tuning likely discovers a more effective non-linear decision logic, enhancing the model’s computational pathway without explicitly improving the linear structure of its representations. In contrast, the $\mathcal{L}_{\text{combined}}$ objective provides a more principled method for enhancing the linear separability of the final embeddings themselves. Thus, while both can surpass the LSC, they do so via different strategies: one by optimizing the downstream computation, the other by directly structuring the representations for that computation.
A trade-off between representation quality and structural robustness. The second key insight is a trade-off between representation quality and generalization to new prompt formats. Methods that are less invasive or only intervene late in the sequence, such as LoRA with $\mathcal{L}_{\text{NT}}$ and especially postfix tuning, demonstrate strong structural robustness, maintaining high performance on OOD prompts. Conversely, methods that either explicitly restructure representations ($\mathcal{L}_{\text{combined}}$) or introduce learnable tokens at the start of the sequence (standard prompt tuning) can exhibit brittleness. For example, on the HOI dataset, LoRA trained with $\mathcal{L}_{\text{combined}}$ shows a significant performance drop on the OOD prompt, despite its final representations remaining highly separable. This suggests that forcing a specific geometric structure can lead to syntactic overfitting, where the model’s generative pathway learns to depend on the training prompt’s structure, compromising its ability to reason flexibly.
Appendix L Results across HOI splits
Table A.6 provides a detailed breakdown of model performance across the four distinct test splits of the Bongard-HOI dataset.
Table A.6: Performance on HOI - detailed breakdown
@@
Phi Baseline gen. $1.5\pm 6.9 55$$5.0\pm 6.9$ $51.0\pm 6.9$ $51.0\pm 6.9$
Baseline sim. $61.5\pm 6.7$ $60.0\pm 6.8$ $57.5\pm 6.9$ $63.0\pm 6.7$
Prompt tuning gen. $60.5\pm 6.8$ $59.0\pm 6.8$ $52.5\pm 6.9$ $62.0\pm 6.7$
Prompt tuning sim. $61.5\pm 6.7$ $59.0\pm 6.8$ $57.0\pm 6.9$ $62.0\pm 6.7$
Prompt tuning gen. $\mathcal{L}_{\text{sim}}$ $66.5\pm 6.5$ $69.0\pm 6.4$ $62.5\pm 6.7$ $63.5\pm 6.7$
Prompt tuning sim. $\mathcal{L}_{\text{sim}}$ $63.5\pm 6.7$ $73.5\pm 6.1$ $70.5\pm 6.3$ $67.5\pm 6.5$
LoRA gen. $\mathcal{L}_{\text{NT}}$ $79.5\pm 5.6$ $79.5\pm 5.6$ $75.5\pm 6.0$ $80.0\pm 5.5$
LoRA sim. $\mathcal{L}_{\text{NT}}$ $65.0\pm 6.6$ $61.5\pm 6.7$ $1.5\pm 6.7 66$$6.5\pm 6.5$
LoRA gen. $\mathcal{L}_{\text{combined}}$ $7.0\pm 5.8 77$$7.5\pm 5.8$ $81.0\pm 5.4$ $81.5\pm 5.4$
LoRA sim. $\mathcal{L}_{\text{combined}}$ $82.0\pm 5.3$ $81.0\pm 5.4$ $77.5\pm 5.8$ $87.5\pm 4.6$
Pixtral Baseline gen. $60.0\pm 6.8$ $61.5\pm 6.7$ $51.5\pm 6.9$ $58.0\pm 6.8$
Baseline sim. $66.5\pm 6.5$ $73.5\pm 6.1$ $69.5\pm 6.4$ $71.5\pm 6.3$
Prompt tuning gen. $60.5\pm 6.8$ $58.5\pm 6.8$ $56.5\pm 6.9$ $59.0\pm 6.8$
Prompt tuning sim. $63.0\pm 6.7$ $62.0\pm 6.7$ $62.5\pm 6.7$ $63.0\pm 6.7$
Prompt tuning gen. $\mathcal{L}_{\text{sim}}$ $73.0\pm 6.2$ $74.0\pm 6.1$ $64.0\pm 6.7$ $72.0\pm 6.2$
Prompt tuning sim. $\mathcal{L}_{\text{sim}}$ $76.0\pm 5.9$ $73.5\pm 6.1$ $70.0\pm 6.4$ $74.5\pm 6.0$
LoRA gen. $\mathcal{L}_{\text{NT}}$ $77.0\pm 5.8$ $79.5\pm 5.6$ $75.5\pm 6.0$ $80.0\pm 5.5$
LoRA sim. $\mathcal{L}_{\text{NT}}$ $3.0\pm 6.2 77$$7.5\pm 5.8$ $71.0\pm 6.3$ $78.0\pm 5.7$
LoRA gen. $\mathcal{L}_{\text{combined}}$ $78.5\pm 5.7$ $81.5\pm 5.4$ $74.5\pm 6.0$ $84.0\pm 5.1$
LoRA sim. $\mathcal{L}_{\text{combined}}$ $79.5\pm 5.6$ $78.5\pm 5.7$ $74.5\pm 6.0$ $78.5\pm 5.7$
Gemma3 4B Baseline gen. $58.5\pm 6.8$ $56.0\pm 6.9$ $53.0\pm 6.9$ $58.5\pm 6.8$
Baseline sim. $50.0$ $50.0$ $50.0$ $50.0$
Prompt tuning gen. $53.0\pm 6.9$ $56.5\pm 6.9$ $53.0\pm 6.9$ $54.0\pm 6.9$
Prompt tuning sim. $50.0$ $50.0$ $50.0$ $50.0$
Prompt tuning gen. $\mathcal{L}_{\text{sim}}$ $56.5\pm 6.9$ $61.5\pm 6.7$ $59.5\pm 6.8$ $58.0\pm 6.8$
Prompt tuning sim. $\mathcal{L}_{\text{sim}}$ $50.0$ $50.0$ $50.0$ $50.0$
LoRA gen. $\mathcal{L}_{\text{NT}}$ $83.5\pm 5.1$ $84.5\pm 5.0$ $85.0\pm 4.9$ $84.0\pm 5.1$
LoRA sim. $\mathcal{L}_{\text{NT}}$ $50.0$ $50.0$ $50.0$ $50.0$
LoRA gen. $\mathcal{L}_{\text{combined}}$ $83.0\pm 5.2$ $84.0\pm 5.1$ $85.0\pm 4.9$ $85.0\pm 4.9$
LoRA sim. $\mathcal{L}_{\text{combined}}$ $84.0\pm 5.1$ $86.5\pm 4.7$ $82.5\pm 5.3$ $80.0\pm 5.5$
@@
There is no significant accuracy degradation when models are tested on unseen objects, unseen actions, or both simultaneously. In fact, for both Phi and Pixtral, the highest generative accuracy with the combined loss is achieved on the most difficult split (unseen obj. / unseen act.), providing strong evidence that the models are learning the abstract relational concept rather than memorizing training examples.
Appendix M Domain generalization
To rigorously assess whether the reasoning skills learned during fine-tuning are truly generalizable, we conducted two distinct cross-domain evaluations. The first tests for generalization across different Bongard-style tasks (OpenWorld and HOI). The second, more challenging evaluation tests whether the learned skills transfer to a task with a completely different structure: the text-image retrieval task of the Winoground dataset.
M.1 Cross-Bongard generalization
@@
In this evaluation, models were fine-tuned on one Bongard dataset (e.g., OpenWorld) and evaluated on the other (e.g., HOI) to assess their ability to generalize beyond the training domain’s specific concepts and structures. The dataset column in Table A.7 indicates the evaluation dataset (i.e., the model was trained on the other dataset).
@@
Table A.7: Cross-Bongard generalization performance. Models were trained on one dataset and evaluated on the other. bold indicates surpassing of LSC. Superscripts on similarity scores denote a significant dependence (p < 0.05) with the predictions of the corresponding generative method (G).
@@
LoRA (HOI, $\mathcal{L}_{\text{NT}}$) $\mathbf{90.0\pm 2.6}$ $84.4\pm 3.2^{G}$ $78.4\pm 3.6^{G}$
LoRA (HOI, $\mathcal{L}_{\text{combined}}$) $\mathbf{92.8\pm 2.3}$ $84.4\pm 3.2^{G}$ $89.2\pm 2.7^{G}$
@@
LoRA (OpenWorld, $\mathcal{L}_{\text{NT}}$) $64.6\pm 3.3$ $71.9\pm 3.1^{G}$ $60.6\pm 3.4$
LoRA (OpenWorld, $\mathcal{L}_{\text{combined}}$) $67.4\pm 3.2$ $71.9\pm 3.1^{G}$ $67.2\pm 3.3^{G}$
@@
LoRA (HOI, $\mathcal{L}_{\text{NT}}$) $\mathbf{88.8\pm 2.8}$ $76.6\pm 3.7^{G}$ $89.0\pm 2.7^{G}$
LoRA (HOI, $\mathcal{L}_{\text{combined}}$) $\mathbf{83.4\pm 3.3}$ $74.4\pm 3.8^{G}$ $87.4\pm 2.9^{G}$
@@
LoRA (OpenWorld, $\mathcal{L}_{\text{NT}}$) $64.9\pm 3.3$ $61.6\pm 3.4^{G}$ $72.5\pm 3.1^{G}$
LoRA (OpenWorld, $\mathcal{L}_{\text{combined}}$) $\mathbf{71.0\pm 3.1}$ $63.1\pm 3.3^{G}$ $68.4\pm 3.2^{G}$
@@
LoRA (HOI, $\mathcal{L}_{\text{NT}}$) $87.2\pm 2.9$ $89.8\pm 2.7^{G}$ $50.0$
LoRA (HOI, $\mathcal{L}_{\text{combined}}$) $83.4\pm 3.3$ $89.8\pm 2.7$ $83.2\pm 3.3^{G}$